The cost for taking a taxi is $1.80 plus $0.10 per eighth of a mile. Write an expression for the cost of a taxi ride that is n miles long. Evaluate your expression to find the cost of a ride that is 4.5 miles long.
Answer by stanbon(75887)

The cost for taking a taxi is $1.80 plus $0.10 per eighth of a mile. Write an expression for the cost of a taxi ride that is n miles long. Evaluate your expression to find the cost of a ride that is 4.5 miles long.
-------------------
C(n) = 1.8 + 0.10(8n)
------------------------
C(4.5) = 1.8 + 0.10(8*4.5)
-----
C(4.5) = 1.8 + 3.6 = $5.40
---------------------
Cheers,
Stan H.
-----------------